OpenStudy (anonymous):

Determine the solution of the system. y = x + 6 y = 2x

OpenStudy (anonymous):

substitute 2 y value 2x into first to get 2x = x + 6 solve for x you get x =6

OpenStudy (anonymous):

2 y i meant second y value = 2x

OpenStudy (anonymous):

could you possible write out the whole formula for me to understand better?

OpenStudy (anonymous):

its not a formula...oh i see what your asking.. 2x = x +6 subtract x from both sides x = 6

OpenStudy (anonymous):

oh i see now! thank you very much!!

OpenStudy (anonymous):

when they give you two equations equaling something else but the variable is same you can equate the two equations to simplify